OpenStudy (anonymous):

|3-2x|=9 solve A. x =-3 b. x = 6 C. x = -3, x = 6 D. x = 3, x = 6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

3-2x=9 , 3-2x=-9 Subtract 3 to both equations Then divide by -2

OpenStudy (anonymous):

3-2x = 9 or -(3-2x) =9 -2x = 6 or 2x-3 = 9 x = -3 or 2x = 12 x=-3 or x = 6

OpenStudy (anonymous):

So answer C is correct :D